Sachs Brands' defined benefit pension plan specifies annual retirement benefits equal to: 1.2% × service years × final year's salary, payable at the end of each year. Angela Davenport was hired by Sachs at the beginning of 2004 and is expected to retire at the end of 2038 after 35 years' service. Her retirement is expected to span 18 years. Davenport's salary is $80,000 at the end of 2018 and the company's actuary projects her salary to be $230,000 at retirement. The actuary's discount rate is 6%. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) At the beginning of 2019, the pension formula was amended to: 1.35% × Service years × Final year's salary The amendment was made retroactive to apply the increased benefits to prior service years. Required: 1. What is the company's prior service cost at the beginning of 2019 with respect to Davenport after the amendment described above? 2. Since the amendment occurred at the beginning of 2019, amortization of the prior service cost begins in 2019. What is the prior service cost amortization that would be included in pension expense? 3. What is the service cost for 2019 with respect to Davenport? 4. What is the interest cost for 2019 with respect to Davenport? 5. Calculate pension expense for 2019 with respect to Davenport, assuming plan assets attributable to her of $160,000 and a rate of return (actual and expected) of 10%. (For all requirements, do not round intermediate calculations. Round your final answers to nearest whole dollar.) 1. Prior service cost _______ 2. Prior service cost amortization _______ 3. Service cost _______ 4. Interest cost _______ 5. Pension expense _______
Answer #1

Part 1)

The value of prior service cost after amendment is arrived as below:

Projected Benefit Obligation (Without Amendment) = (1.2%*Service Years*Final Year's Salary)*PVA(Discount Rate,Years)*PV(Discount Rate,Years) = (1.2%*15*230,000)*PVIFA(6%,18)*PVIF(6%,20) = 41,400*10.82760*.31180 = $139,768

Projected Benefit Obligation (With Amendment) = (1.35%*Service Years*Final Year's Salary)*PVA(Discount Rate,Years)*PV(Discount Rate,Years) = (1.35%*15*230,000)*PVIFA(6%,18)*PVIF(6%,20) = 46,575*10.82760*.31180 = $157,239

Now, we can determine prior service cost as below:

Prior Service Cost After Amendment = Projected Benefit Obligation (With Amendment) - Projected Benefit Obligation (Without Amendment) = 157,239 - 139,768 = $17,471

______

Part 2)

The value of prior service cost amortization is determined as below:

Prior Service Cost Amortization Included in Pension Expense = Prior Service Cost After Amendment/Expected Remaining Service = 17,471/20 = $874

______

Part 3)

The service cost for 2019 is arrive as follows:

Service Cost for 2019 = (1.35%*Service Years*Final Year's Salary)*PVA(Discount Rate,Years)*PV(Discount Rate,Years) = (1.35%*1*230,000)*PVA(6%,18)*PV(6%,19) = 3,105*10.82760*0.33051 = $11,112

______

Part 4)

The amount of interest cost for 2019 is calculated as below:

Interest Cost for 2019 = Projected Benefit Obligation (With Amendment)*Discount Rate = 157,239*6% = $9,434

______

Part 5)

The value of pension expense for 2019 is determined as follows:

Pension Expense for 2019 = Service Cost for 2019 + Interest Cost for 2019 - Return on Plan Assets + Amortization of Prior Service Cost = 11,112 + 9,434 - 160,000*10% + 874 = $5,420

______

Notes:

1) There can be a slight difference in final answers on account of rounding off values.

2) PVA indicates Present Value of Annuity of $1 and PV indicates Present Value of $1. The values have been derived with the use of Present Value tables.